# State Space Model Meets Transformer: A New Paradigm for 3D Object Detection

### v1

@@ 2 Related Work
In contrast, indoor 3D object detection involves handling a more diverse set of object categories and shapes, as well as more complex spatial relationships between objects. Existing indoor 3D object detection methods can be broadly categorized into three groups: vote-based methods, expansion-based methods, and DETR-based methods. For vote-based methods, VoteNet (Qi et al., 2019), as the pioneering work, designs a novel 3D proposal mechanism based on deep Hough voting. MLCVNet (Xie et al., 2020) introduces three context modules into the voting and classifying stages of VoteNet to encode contextual information at different levels. BRNet (Cheng et al., 2021) back-traces the representative points from the vote centers to better capture the fine local structural features surrounding the potential objects from the raw point clouds. H3DNet (Zhang et al., 2020) predicts a hybrid set of geometric primitives and converts the predicted geometric primitives into object proposals. For expansion-based methods, GSDN (Gwak et al., 2020) proposes a generative sparse tensor decoder to generate virtual center features from surface features while discarding unlikely object centers. FCAF3D (Rukhovich et al., 2022) further introduces a fully convolutional anchor-free indoor 3D object detection method. CAGroup3D (Wang et al., 2022a) generates high-quality 3D proposals by leveraging a class-aware local grouping strategy on object surface voxels with consistent semantic predictions. The above methods require carefully designed proposal generation modules and involve several manually set thresholds. DETR (Carion et al., 2020) is a pioneering work that applies Transformers (Vaswani, 2017) to 2D object detection, eliminating many hand-crafted components such as Non-Maximum Suppression (Neubeck & Van Gool, 2006) or anchor boxes (Girshick, 2015; Ren, 2015; Lin, 2017). Currently, DETR and its variants (Zhu et al., 2020; Dai et al., 2021; Liu et al., 2022) have achieved state-of-the-art results in various 2D object detection tasks. Inspired by these works, numerous DETR-based 3D object detection methods have been explored. 3DETR (Misra et al., 2021) is the first to introduce an end-to-end Transformer model for 3D object detection, achieving promising results. GroupFree (Liu et al., 2021) employs a key point sampling strategy to select candidate points and utilizes the attention mechanism to update query point features. LeadNet (Wang et al., 2023) further improves the transformer decoder by introducing a dynamic object query sampling module and a dynamic Gaussian weight map. Most recently, VDETR (Shen et al., 2024) proposes a novel 3D vertex relative position encoding method, which directs the model to focus on points near the object, achieving state-of-the-art performance. However, these DETR-based methods use fixed scene point features in different decoder layers, which limits the detection capabilities of later layers. Unlike the above methods, we propose an ISSM-based decoder that simultaneously updates both scene point and query point features.
@@ 3.4 ISSM-based Decoder
IBS module. To facilitate bidirectional interaction among different scene points in a single pass, we follow Vision Mamba (Zhu et al., 2024) and introduce a bidirectional scanning mechanism into our ISSM. Figure 4 (b) illustrates the detailed structure of the IBS module. Firstly, we input the forward-ordered and backward-ordered scene points into their respective forward and backward ISSMs. Both the forward and backward ISSMs use state points as system states, with each ISSM generating the updated features of the scene points and state points as outputs. We then fuse the outputs using a linear layer to obtain the final features of the scene points and state points. Additionally, we incorporate a depthwise convolution (Chollet, 2017) for local feature extraction of the scene points. In Appendix A.2, we provide detailed algorithmic procedures of the IBS module.
GFFN. The gating mechanism (Dauphin et al., 2017) introduces gated linear units, allowing the model to dynamically select activation paths based on the input. This selectivity enhances the model flexibility, enabling the model to better capture complex patterns. Thus, we design the GFFN to replace the standard FFN, as shown in Figure 4 (c). For the ordered scene points, we also incorporate a depthwise convolution to fully leverage the spatial structural information of the point cloud.
@@ 3.5 Model Setups
For a fair comparison, we build our DEST-based 3D detector based on two baseline DETR-based models: GroupFree (Liu et al., 2021) and VDETR (Shen et al., 2024), respectively. For GroupFree baseline, the input points $P_{in}\in\mathbb{R}^{N\times 3}$ contain only position information. We replace all transformer decoder layers with our proposed ISSM-based decoder layers while retaining the original spatial encodings and the original detection heads. For the training loss, we employ the same detection loss with GroupFree and introduce an additional objectness loss for the scene points. Specifically, we use a MLP head to determine whether a scene point is a foreground point and apply a binary focal loss to supervise the prediction results. For VDETR baseline, the input points $P_{in}\in\mathbb{R}^{N\times 6}$ include both position and RGB information. In the decoder, we also replace all transformer decoder layers with our ISSM-based decoder layers while retaining the original detection heads. Regarding the spatial encodings, we generate them for the corresponding state points using the predicted 3D bounding boxes and for the scene points using their point positions. For the training loss, we employ the same detection loss with VDETR and introduce the same objectness loss as above for the scene points. Due to space limitations, more details can be found in the Appendix.
@@ 4.3 Ablation Experiments
In this section, we verify the key design moduels of our DEST-based detector. All ablation experiments are conducted on the ScanNet V2 dataset with the GroupFree(S) baseline. Following GroupFree (Liu et al., 2021), we report the average performance of 25 trials by default.
@@ A.4.1 Model Setup Details
GroupFree baseline. Given the input points $P_{in}\in\mathbb{R}^{N\times 3}$ without RGB information, we employ PointNet++ (Qi et al., 2017b) as the point cloud encoder for a fair comparison. In addition, we utilize the same initial object candidate sampling module as employed in GroupFree (Liu et al., 2021). However, unlike GroupFree, where the object candidate points are used as the initial query points, they are treated as the initial state points in our ISSM-based model. In the decoder, we replace all transformer decoder layers with our proposed ISSM-based decoder layers while retaining the original spatial encodings and the original detection heads.
VDETR baseline. Given the input points $P_{in}\in\mathbb{R}^{N\times 6}$ with RGB information, we employ the same 3D sparse convolution network of VDETR (Shen et al., 2024) as the point cloud encoder for a fair comparison. Additionally, we utilize the same initial object query sampling module as employed in VDETR (Shen et al., 2024) to sample the initial state points. In the decoder, we also replace all transformer decoder layers with our proposed ISSM-based decoder layers while retaining the original detection heads. Regarding the spatial encodings, we generate them for the corresponding state points using the predicted 3D bounding boxes and for the scene points using their point positions.
@@ A.6 Visual Comparison
In this section, we present a visual comparison of the prediction results of our method and DETR-based methods. To ensure a fair comparison, we apply the same post-processing steps to the detection results. First, we filter out low-quality predictions using a class confidence threshold of 0.3, followed by class-specific Non-Maximum Suppression with an IoU threshold of 0.5 to remove redundant bounding boxes. As shown in Figure 6, Figure 7, and Figure 8, we first compare the prediction results of our method with three baseline methods on the ScanNet dataset. Our method successfully detects most target objects across various scenes, with the predicted bounding boxes being more closely aligned with the ground truth labels. Subsequently, we present a visual comparison of the prediction results on the SUN RGB-D dataset. Since VDETR does not provide an official implementation for the SUN RGB-D dataset, we only include a visual comparison with the GroupFree baseline as shown in Figure 9. Since VDETR does not provide an official implementation for the SUN RGB-D dataset, we only include a visual comparison with the GroupFree baseline. Unlike the ScanNet dataset, the SUN RGB-D dataset is generated from single images, which results in point cloud data with more pronounced occlusions and uneven density, making the detection task more challenging. Nevertheless, our method still achieves visually satisfactory detection results.

### v2

@@ 2 Related Work
In contrast, indoor 3D object detection involves handling a more diverse set of object categories and shapes, as well as more complex spatial relationships between objects. Existing indoor 3D object detection methods can be broadly categorized into three groups: vote-based methods, expansion-based methods, and DETR-based methods. For vote-based methods, VoteNet (Qi et al., 2019), as a pioneering work, designs a novel 3D proposal mechanism based on deep Hough voting. MLCVNet (Xie et al., 2020) introduces three context modules in the voting and classifying stages of VoteNet to encode contextual information at different levels. BRNet (Cheng et al., 2021) backtraces representative points from the voting centers to better capture the fine local structural features surrounding the potential objects from the raw point clouds. H3DNet (Zhang et al., 2020) predicts a hybrid set of geometric primitives and converts the predicted geometric primitives into object proposals. For expansion-based methods, GSDN (Gwak et al., 2020) proposes a generative sparse tensor decoder to generate virtual center features from surface features while discarding unlikely object centers. FCAF3D (Rukhovich et al., 2022) further introduces a fully convolutional anchor-free indoor 3D object detection method. CAGroup3D (Wang et al., 2022a) generates high-quality 3D proposals by leveraging a class-aware local grouping strategy on object surface voxels with consistent semantic predictions. The above methods require carefully designed proposal generation modules and involve several manually set thresholds. DETR (Carion et al., 2020) is a pioneering work that applies Transformers (Vaswani, 2017) to 2D object detection, eliminating many hand-crafted components such as Non-Maximum Suppression (Neubeck & Van Gool, 2006) or anchor boxes (Girshick, 2015; Ren, 2015; Lin, 2017). Currently, DETR and its variants (Zhu et al., 2020; Dai et al., 2021; Liu et al., 2022) have achieved state-of-the-art results in various 2D object detection tasks. Inspired by these works, numerous DETR-based 3D object detection methods have been explored. 3DETR (Misra et al., 2021) is the first to introduce an end-to-end Transformer model for 3D object detection, achieving promising results. GroupFree (Liu et al., 2021) employs a key point sampling strategy to select candidate points and utilizes the attention mechanism to update query point features. LeadNet (Wang et al., 2023) further improves the transformer decoder by introducing a dynamic object query sampling module and a dynamic Gaussian weight map. Most recently, VDETR (Shen et al., 2024) proposes a novel 3D vertex relative position encoding method, which directs the model to focus on points near the object, achieving state-of-the-art performance. However, these DETR-based methods use fixed scene point features in different decoder layers, which limits the detection capabilities of later layers. Unlike the above methods, we propose an ISSM-based decoder that simultaneously updates both scene point and query point features.
@@ 3.4 ISSM-based Decoder
IBS module. To facilitate bidirectional interaction among different scene points in a single pass, we follow Vision Mamba (Zhu et al., 2024) and introduce a bidirectional scanning mechanism into our ISSM. Figure 4 (b) illustrates the detailed structure of the IBS module. Firstly, we input the forward-ordered and backward-ordered scene points into their respective forward and backward ISSMs. Both the forward and backward ISSMs use state points as system states, with each ISSM generating the updated features of the scene points and state points as outputs. We then fuse the output using a linear layer to obtain the final features of the scene points and state points. Additionally, we incorporate a depthwise convolution (Chollet, 2017) for local feature extraction of the scene points. In Appendix A.2, we provide detailed algorithmic procedures of the IBS module.
GFFN. The gating mechanism (Dauphin et al., 2017) introduces gated linear units, which allows the model to dynamically select activation paths based on the input. This selectivity improves the flexibility of the model, enabling the model to better capture complex patterns. Thus, we design the GFFN to replace the standard FFN, as shown in Figure 4 (c). For the ordered scene points, we also incorporate a depth-wise convolution to fully leverage the spatial structural information of the point cloud.
@@ 3.5 Model Setups
For a fair comparison, we build our DEST-based 3D detector based on two baseline DETR-based models: GroupFree (Liu et al., 2021) and VDETR (Shen et al., 2024), respectively. For GroupFree baseline, the input points $P_{in}\in\mathbb{R}^{N\times 3}$ contain only position information. We replace all transformer decoder layers with our proposed ISSM-based decoder layers while retaining the original spatial encodings and the original detection heads. For the training loss, we employ the same detection loss with GroupFree and introduce an additional objectness loss for the scene points. Specifically, we use an MLP head to determine whether a scene point is a foreground point and apply a binary focal loss to supervise the prediction results. For VDETR baseline, the input points $P_{in}\in\mathbb{R}^{N\times 6}$ include both position and RGB information. In the decoder, we also replace all transformer decoder layers with our ISSM-based decoder layers while retaining the original detection heads. Regarding the spatial encoding, we generate them for the corresponding state points using the predicted 3D bounding boxes and for the scene points using their point positions. For the training loss, we employ the same detection loss with VDETR and introduce the same objectness loss as above for the scene points. Due to space limitations, more details can be found in the Appendix.
@@ 4.3 Ablation Experiments
In this section, we verify the key design modules of our DEST-based detector. All ablation experiments are conducted on the ScanNet V2 dataset with the GroupFree(S) baseline. Following GroupFree (Liu et al., 2021), we report the average performance of 25 trials by default.
@@ A.4.1 Model Setup Details
GroupFree baseline. Given the input points $P_{in}\in\mathbb{R}^{N\times 3}$ without RGB information, we employ PointNet++ (Qi et al., 2017b) as the point cloud encoder for a fair comparison. In addition, we utilize the same initial object candidate sampling module as employed in GroupFree (Liu et al., 2021). However, unlike GroupFree, where the object candidate points are used as the initial query points, they are treated as the initial state points in our ISSM-based model. In the decoder, we replace all transformer decoder layers with our proposed ISSM-based decoder layers while retaining the original spatial encoding and the original detection heads.
VDETR baseline. Given the input points $P_{in}\in\mathbb{R}^{N\times 6}$ with RGB information, we employ the same 3D sparse convolution network of VDETR (Shen et al., 2024) as the point cloud encoder for a fair comparison. Additionally, we utilize the same initial object query sampling module as employed in VDETR (Shen et al., 2024) to sample the initial state points. In the decoder, we also replace all transformer decoder layers with our proposed ISSM-based decoder layers while retaining the original detection heads. Regarding the spatial encoding, we generate them for the corresponding state points using the predicted 3D bounding boxes and for the scene points using their point positions.
@@ A.6 Visual Comparison
In this section, we present a visual comparison of the prediction results of our method and DETR-based methods. To ensure a fair comparison, we apply the same post-processing steps to the detection results. First, we filter out low-quality predictions using a class confidence threshold of 0.3, followed by class-specific Non-Maximum Suppression with an IoU threshold of 0.5 to remove redundant bounding boxes. As shown in Figure 6, Figure 7, and Figure 8, we first compare the prediction results of our method with three baseline methods on the ScanNet V2 dataset. Our method successfully detects most target objects across various scenes, with the predicted bounding boxes being more closely aligned with the ground truth labels. Subsequently, we present a visual comparison of the prediction results on the SUN RGB-D dataset. Since VDETR does not provide an official implementation for the SUN RGB-D dataset, we only include a visual comparison with the GroupFree baseline as shown in Figure 9. Since VDETR does not provide an official implementation for the SUN RGB-D dataset, we only include a visual comparison with the GroupFree baseline. Unlike the ScanNet dataset, the SUN RGB-D dataset is generated from single images, which results in point cloud data with more pronounced occlusions and uneven density, making the detection task more challenging. Nevertheless, our method still achieves visually satisfactory detection results.